# Rajamanickam

### v1

@@ 2 Remarks on the strong anchoring condition in three dimensions
$\mathbf{Q}_{homo}=\begin{cases}\mathbf{0}&\quad\text{for }A>B^{2}/27C,\\ s_{+}(\mathbf{n}_{+}\otimes\mathbf{n}_{+}-\tfrac{1}{2}\mathbf{I})&\quad\text{for }A<B^{2}/27C\end{cases}$ (2)
@@
which determines the distance over which nematic order is locally suppressed and thus sets the size of defect cores. At the nematic–isotropic transition, $\xi=\sqrt{27LC/B^{2}}$, while deep in the nematic phase, $\xi\approx\sqrt{L/|A|}$.
By comparing the two length scales introduced above, we can infer that
  • $l_{ex}\ll\xi$: Surface anchoring dominates down to distances of order $l_{ex}$ and the boundary defects of size $\xi$ are strongly influenced by the surface anchoring. In other words, Dirichlet boundary effectively hold even within the defects.
@@ 3 Problem formulation in the reduced two-dimensional Landau–de Gennes framework
where $A=A(T)$ is a linear function of the temperature $T$, while $L$ and $C$ are positive material constants; the anchoring strength $W$ will be assumed positive and $\mathbf{Q}_{b}$ is the preferred value of $\mathbf{Q}$ on $\partial\Omega$. The isotropic phase corresponds to $\mathbf{Q}=0$, whereas, the nematic phase with orientational order corresponds to $\mathbf{Q}\neq 0$. In the reduced model, $\mathrm{tr}\,\mathbf{Q}^{3}=0$, which rules out biaxiality, and the reduced $\mathbf{Q}$-tensor can be written as
@@
is the ratio of the domain size $h$ to the 2D reduced coherence length $\xi=\sqrt{L/|A|}$. The surface-energy term now implies that, as the domain size becomes large ($\varepsilon\to 0$), it grows proportionally indicating a strong anchoring boundary condition, although in very small domains ($\varepsilon\to\infty$), it must become weak. The latter implication is expected since the domain has simply shrunk to the scale of a defect core radius.
The Euler–Lagrange equations for $q_{1}$ and $q_{2}$ are given by
@@ 5 Oseen–Frank limit, $\varepsilon\to\infty$
Circle 11 one +1+1-interior defect
@@
Below, we provide a prescription of the solution in the asymptotic limit $\varepsilon\to\infty$ using the language of method of matched asymptotic expansions. The description This construction applies to any admissible solution of the leading-order problem.
@@ 6 Numerical results with tangential anchoring
In this section, we present numerical solutions of the problem (12)-(14) for few illustrative domains subjected to tangential anchoring, with $s_{b}=2\sqrt{q_{1,b}^{2}+q_{2,b}^{2}}=1$. The stability of the computed solutions is also evaluated; the details of the stability analysis are provided in Appendix A Appendix A: Formulation of the stability analysis. In all numerical computations, we set $\gamma=1$. For comparison, we also consider the $\varepsilon$-independent Dirichlet boundary condition
$q_{1}=q_{1,b},\quad q_{2}=q_{2,b}\quad\text{on}\quad\partial\Omega.$ (68)

### v2

@@ 2 Remarks on the strong anchoring condition in three dimensions
$\mathbf{Q}_{homo}=\begin{cases}\mathbf{0}&\quad\text{for }A>B^{2}/27C,\\ s_{+}(\mathbf{n}_{+}\otimes\mathbf{n}_{+}-\tfrac{1}{3}\mathbf{I})&\quad\text{for }A<B^{2}/27C\end{cases}$ (2)
@@
which determines the distance over which nematic order is locally suppressed and thus sets the size of defect cores. At the nematic–isotropic transition, $\xi=\sqrt{27LC/B^{2}}$, while deep in the nematic phase, $\xi\approx\sqrt{L/2|A|}$. By comparing the two length scales introduced above, we can infer that
  • $l_{ex}\ll\xi$: Surface anchoring dominates down to distances of order $l_{ex}$ and the boundary defects of size $\xi$ are strongly influenced by the surface anchoring. In other words, Dirichlet boundary conditions effectively hold even within the defects.
@@ 3 Problem formulation in the reduced two-dimensional Landau–de Gennes framework
In the reduced model, $\mathrm{tr}\,\mathbf{Q}^{3}=0$, which rules out biaxiality, and the reduced $\mathbf{Q}$-tensor can be written as
@@
is the ratio of the domain size $h$ to the 2D reduced coherence length $\xi=\sqrt{L/|A|}$. The surface-energy term now implies that, as the domain size becomes large ($\varepsilon\to\infty$), it grows proportionally indicating a strong anchoring boundary condition, although in very small domains ($\varepsilon\to 0$), it must become weak. The Euler–Lagrange equations for $q_{1}$ and $q_{2}$ are given by
@@ 5 Oseen–Frank limit, $\varepsilon\to\infty$
Circle 22 Two $\tfrac{1}{2}-$interior defects
@@
Below, we provide a prescription of the solution in the asymptotic limit $\varepsilon\to\infty$ using the language of method of matched asymptotic expansions. This construction applies to any admissible solution of the leading-order problem.
@@ 6 Numerical results with tangential anchoring
In this section, we present numerical solutions of the problem (12)-(14) for few illustrative domains subjected to tangential anchoring, with $s_{b}=2(q_{1,b}^{2}+q_{2,b}^{2})^{\frac{1}{2}}=1$. The stability of the computed solutions is also evaluated; the details of the stability analysis are provided in Appendix A. In all numerical computations, we set $\gamma=1$. For comparison, we also consider the $\varepsilon$-independent Dirichlet boundary condition
$q_{1}=q_{1,b},\qquad q_{2}=q_{2,b}\quad\text{on}\quad\partial\Omega.$ (68)